OpenStudy (anonymous):

solve x^3 + 20x = -2x^3 - 7x 3x^3 + 27x = 0 3x (x^2 + 9) = 0 3x (x - 3) (x + 3) = 0 x = {-3, 0, 3} describe and correct the error by step please :)

OpenStudy (anonymous):

there is a miss take..... 3x(x^2+9)=0 ...... goood but 3x(x-3)(x+3)=0 is wrong

OpenStudy (anonymous):

it should be 3x(x^2+3^2)=0

OpenStudy (anonymous):

your final answer is right

OpenStudy (anonymous):

how can someone solve it exactly if it were another type of problem?

OpenStudy (anonymous):

x^2+9=0

OpenStudy (anonymous):

x^2=-9 x=3i

OpenStudy (anonymous):

x=+- 3i

OpenStudy (anonymous):

3x (x - 3) (x + 3) = 0 3x[x^2+3x-3x-9]=0 3x(x^2-9)=0 but you have 3x(x^2+9)

OpenStudy (anonymous):

you just saved my life! xD thank youuuuu